40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7949     Entry time: Mon Jan 28 21:32:38 2013
Author: jamie 
Type: Update 
Category: Alignment 
Subject: tweaking of alignment into half PRC 

[Koji, Jamie]

We tweaked up the alignment of the half PRC a bit.  Koji started by looking at the REFL and POP DC powers as a function of TT2 and PRM alignment. 
He found that the reflected beam for good PRC transmission was not well overlapped at REFL.  When the beam was well overlapped at REFL, there was clipping in the REFL path on the AS table.

We started by getting good overlap at REFL, and then went to the AS table to tweak up all the beams on the REFL pds and cameras.
This made the unlocked REFL DC about 40 count. This was about 10mV (=0.2mA) at the REFL55 PD.
This amazed Koji since we found the REFL DC (of the day) of 160 as the maximum of the day for a particular combination of the PRM Pitch and TT2 Pitch. So something wrong could be somewhere.

We then moved to the ITMX table where we cleaned up the POP path.  We noticed that the lens in the POP path is a little slow, so the beam is too big on the POP PD and on the POP camera (and on the camera pick-off mirror as well)
We moved the currently unused POP55 and POP22/110 RFPDs out of the way so we could move the POP RF PD and camera back closer to the focus.  Things are better, but we still need to get a better focus, particularly on the POP PD.

We found two irides on the oplev path. They are too big and one of these is too close to the POP beam. Since it does not make sense too to have two irides in vicinity, we pulled out that one from the post.

Other things we noticed:

  • The POP beam is definitely clipping in the vacuum, looks like on two sides.
  • We can probably get better layout on the POP table, so we're not hitting mirrors at oblique angles and can get beams on grid paths.

After the alignment work on the tables, we started locking the cavity. We already saw the improvement of the POPDC power from 1000 cnt to 2500 cnt without any realignment.
Once PRM is tweaked a little (0.01ish for pitch and yaw), the maximum POPDC of 6000 was achieved. But still the POP camera shows non-gaussian shape of the beam and the Faraday camera shows bright
scattering of the beam. It seems that the scattering at the Faraday is not from the main beam but the halo leaking from the cavity (i.e. unlocking of the cavity made the scattering disappeared)


Tomorrow Jenne and I will go into BS to tweak the alignment of the TEMP PRC flat mirror, and into ITMX to see if we can clean up the POP path.

ELOG V3.1.3-